OpenStudy (anonymous):

Which fraction is between 5/10 and 1

OpenStudy (anonymous):

some 1 :( i need help badly

OpenStudy (anonymous):

Just use the midpoint formula (x1 + x2)/2 = midpoint

OpenStudy (anonymous):

Thanks daltino, and okbuw i dont get it ure the best

OpenStudy (anonymous):

wait so whats the answer

OpenStudy (anonymous):

Anything between 1/2 and 1

OpenStudy (anonymous):

ok thanks daltino and okubow

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

Who ever helps me ill prank call some one with prankdial.com from credit fraud lol

OpenStudy (anonymous):

(5/10 + 1)/2 (5/10 + 10/10)/2 (15/10)/2 15/20 3/4

OpenStudy (anonymous):

Oh yeah, I thought you meant any fraction between the two, but 3/4 is the one in the dead center of 1/2 and 1, he is right.

OpenStudy (anonymous):

yeah i ned LCD FOR both fractions

OpenStudy (anonymous):

5/18 and 3/16

OpenStudy (anonymous):

This question is just testing your understanding of the midpoint formula. They want to see if you know the formula and know when to apply it. All you need to do is to add the larger number to the smaller number and then divide by 2. That's the midpoint formula.

OpenStudy (anonymous):

upps sorry wrong post i mean 5/18 and 3/16 i need the LCD (Least Common Denominator)

OpenStudy (anonymous):

5/18= 80/288(reduce that) 3/16= 26/288(also reduce)

OpenStudy (anonymous):

Its a glitch i never askedwich fraction whas beetween

OpenStudy (anonymous):

54/288 rather

OpenStudy (anonymous):

so answer is 80/288 and 26/288?

OpenStudy (anonymous):

no, 54/288, not 26/288.

OpenStudy (anonymous):

But remember to reduce.

OpenStudy (anonymous):

upps i forgot to reduce lol and i have a pen

OpenStudy (anonymous):

27/144 13/144

OpenStudy (anonymous):

dude ure the best thank you so much

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

u online everyday lol

OpenStudy (anonymous):

No, I signed up like an hour ago haha.

OpenStudy (anonymous):

okubow thank you so much i had really appriciated yall 2

OpenStudy (anonymous):

daltino log in tommorow lol ure my math teahcer for free answers lol